### v1

# Enhanced Expressivity in Graph Neural Networks with Lanczos-Based Linear Constraints

Abstract
[…]
Introduction
Graphs play a crucial role in various domains, representing linked data such as social networks [1], citation networks [42], knowledge graphs [35], metabolic network reconstruction [37], and user-item graphs in recommender systems [32]. Graph Neural Networks (GNNs) have emerged as state-of-the-art tools for processing graph-structured data. Message Passing Neural Networks (MPNNs) is the most prevalent technique within GNNs, which, relying on neighborhood aggregation, exhibit expressiveness no greater than the first-order Weisfeiler-Leman (1-WL) test [46, 34, 47, 33]. Therefore, MPNNs cannot distinguish specific graph structures, e.g., $k$-regular graphs. Moreover, link prediction (LP) tasks cannot always be answered reliably based on pairs of node embeddings obtained from MPNNs. Specifically, the node automorphism problem arises in instances where two nodes possess identical local structures, resulting in equivalent embeddings and, consequently, identical predictions. However, their relationships to a specific node, e.g., in terms of distance, may differ [55, 13].
Efforts aimed at augmenting the expressiveness of MPNNs have pursued four main directions: aligning with the $k$-WL hierarchy [34, 31, 4], enriching node features with identifiers, exploiting structural information that cannot be captured by the WL test [9, 8], and Subgraph GNNs (SGNNs) [6, 20]. SGNNs are a recent class of expressive GNNs that model graphs through a collection of subgraphs. They have emerged as a potential solution, extending GNNs by collecting extracted subgraphs explicitly or implicitly and incorporating these into the GNN architecture. Subgraph extraction can be achieved, e.g., by removing or marking specific nodes or directly counting specific substructures [38]. However, in the worst case, previous SGNNs involve computationally intensive preprocessing steps or running a GNN many times.
To address these limitations, we introduce a novel approach grounded in graph signal processing [36] and spectral graph theory [15]. Our method introduces a novel eigenbasis, denoted as the Learnable Lanczos with Linear Constraints (LLwLC), which can explicitly encode linear constraints, particularly those derived from extracted induced subgraphs, into the basis. We propose a low-rank approximation [16] of the Laplacian matrix based on the Lanczos algorithm with linear constraints [18].
[…]
Figure 1: The comprehensive workflow of LLwLC: We introduce a novel eigenbasis to boost the expressivity of GNNs by satisfying constraints $\mathbf{C}$. For the two $2$-regular input graphs (first dotted block), the constraint matrix $\mathbf{C}$ (two columns shown in the second dotted block) is derived. Specifically, as depicted, for the first Neumann eigenvalue constraint (focused on link representation), we integrate the degrees of nodes into the constraint matrix such that the summation over adjacent nodes in the boundary $x\sim y$ ensures $\bm{f}(x)-\bm{f}(y)=0$. To predict the dotted link, nodes up to two hops away are analyzed: one-hop in light blue and two-hop in dark blue, centered on yellow query nodes. Based on the input graph Laplacian matrix $\mathbf{L}$ and the constraint matrix $\mathbf{C}$, we address the eigenvalue problem under these linear constraints using the Lanczos algorithm with the linear constraint consisting of an outer loop (Lanczos algorithm) and an inner loop ($\mathbf{P}(\nu)$ to $\mathcal{N}(\mathbf{C})$ with LSQR algorithm) resulting in the eigenpairs $\mathbf{V}=\mathbf{Q}\mathbf{B}$ and $\mathbf{R}$. The third block demonstrates constructing eigenbases for the two 2-regular graphs, ensuring features conform to constraints (indicated in blue). We stack multiple blocks, where each learns the new features by applying a multi-layer perceptron ($f$) over the eigenvalue matrix $\mathbf{R}$, so we have $\sigma(\mathbf{V}f(\mathbf{R})\mathbf{V}{{}^{\top}}\mathbf{X}\mathbf{W})$. This sequence concludes with global pooling and a fully connected layer to output the predicted link probability $\hat{y}$.
The LLwLC eigenbasis enhances feature expressiveness by incorporating linear constraints derived from the graph structure. We propose two novel subgraph extraction policies, focusing on vertex-deleted subgraphs and Neumann eigenvalue constraints. We provide a conjecture that vertex-deleted subgraphs have the universal approximator capabilities of LLwLC, while Neumann constraints enable the encoding of boundary conditions and link representations between nodes, allowing to differentiate $k$-regular graphs that are indistinguishable by WL. Theoretical analysis indicates that LLwLC can be applied in various problem settings. We evaluate its effectiveness in link prediction tasks, where expressivity and particularly the ability to distinguish automorphic nodes is pivotal.
Our research presents several key contributions: (i) We develop a novel, efficient eigenbasis (LLwLC) for encoding linear constraints, including induced subgraphs, using the Lanczos algorithm with linear constraints. (ii) We explore the application of Neumann eigenvalue constraints within this eigenbasis, which encodes induced subgraphs and link representations. (iii) We provide a comprehensive theoretical analysis of LLwLC’s convergence. (iv) We investigate the impact of vertex-deleted subgraphs on improving expressiveness of our proposed LLwLC. (v) We conduct extensive experiments showcasing the significant impact of the LLwLC on the challenging link prediction task.
Preliminaries
Notations
A graph $G(V,E,\mathbf{X})$ consists of a vertex set $V$, edge set $E$, and node features $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $n$ is the number of nodes. Each column $\bm{x}_{v}\in\mathbb{R}^{d}$ represents the features of node $v\in V$. $\mathbf{A}$ and $\mathbf{D}$ are the graph’s adjacency and degree matrices, respectively. The graph Laplacian is $\mathbf{L}=\mathbf{D}-\mathbf{A}$, with $\mathbf{U}$ and $\mathbf{\Lambda}$ as its eigenvector and eigenvalue matrices. The degree of node $v$ is denoted by $d_{v}$.
Spectral Graph Convolutional Networks
For a graph signal $\bm{x}\in\mathbb{R}^{n}$[43] defines the graph Fourier transform, $\mathbf{U}^{\top}\bm{x}$, and its inverse, $\mathbf{U}\bm{x}$, based on the eigenbasis of the graph Laplacian matrix $\mathbf{L}$. The graph convolution is $\mathbf{U}(\mathbf{U}^{\top}\bm{x}\ast\mathbf{U}^{\top}\bm{y})=\mathbf{U}g(\mathbf{\Lambda})\mathbf{U}^{\top}\bm{x}$ where $\bm{y}$ is the graph filter, $g$ is the function applied over the eigenvalue matrix $\mathbf{\Lambda}$ to encode the graph filter, and $\ast$ is the elementwise multiplication. The seminal spectral GCN method [12] is cubic in the number of nodes. To address this, different $g$ functions were defined [23, 3]. LanczosNet [29] uses the Lanczos algorithm for fast multi-scale computation with learnable spectral filters. However, like previous GNNs, LanczosNet has limited expressivity, e.g., it cannot distinguish between the $k$-regular graphs. To address these limitations and enhance feature expressiveness in GNNs, we introduce a novel learnable spectral basis for encoding subgraphs as linear constraints.
Lanczos with Linear Constraint Networks
This section outlines the Lanczos Algorithm with Linear Constraints and the construction of a constraint matrix $\mathbf{C}$ for subgraph embedding, followed by the development of the complete LLwLC block and LLwLCNet pipeline. It also includes a proof of convergence properties and explores subgraph extraction policies, focusing on Neumann eigenvalue constraints and vertex-deleted subgraphs.
Algorithm 1 Lanczos Algorithm with Linear Constraint (LLwLC). Main steps for subgraph encoding in the Graph Laplacian eigenbasis are highlighted in red.
1:Input: $\mathbf{L},\mathbf{P}=\mathbf{I}-\mathbf{C}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}$, random initialized vector $\bm{\nu},$ $\kappa-$step LLwLC algorithm, error tolerance $\epsilon$
2:Init: $\bm{\nu}_{1}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{P}}(\bm{\nu}),\beta_{1}=\lVert\bm{\nu}_{1}\rVert_{2},\bm{q}_{0}=0$
3:for $j=1$ to $\kappa$ do
4:     $\bm{q}_{j}=\frac{\bm{\nu}_{j}}{\beta_{j}}$
5:     $\bm{p}_{j}=\mathbf{P}(\mathbf{L}\bm{q}_{j})$
6:     $\bm{u}_{j}=\mathbf{L}\bm{q}_{j}-\beta_{j}\bm{q}_{j-1}$rightarrow $\bm{p}_{j}-\beta_{j}\bm{q}_{j-1}$
7:     $\alpha_{j}=\bm{u}_{j}^{\top}\bm{q}_{j}$
8:     $\bm{\nu}_{j+1}=\bm{u}_{j}-\alpha_{j}\bm{q}_{j}$
9:     $\beta_{j+1}=\lVert\bm{\nu}_{j+1}\rVert_{2}$
10:     if $\beta_{j+1}\leq\epsilon$ then
11:         quit
12:     end if
13:end for
14:$Q=[q_{1},\dots,q_{\kappa}]$, Construct $\mathbf{T}$
15:$\text{EVD}(\mathbf{T})=\mathbf{B}\mathbf{R}\mathbf{B}^{\top}$
16:return $\mathbf{V}=\mathbf{Q}\cdot\mathbf{B}$ and $\mathbf{R}$
Algorithm 2 LLwLCNet
[…]
2:Init: $Y_{0}=X$
3:for $j=0$ to $k-1$ do
4:     $\mathbf{Y}_{i+1}=\sigma(\mathbf{V}f_{i}(\mathbf{R})\mathbf{V}^{\top}\mathbf{Y}_{i}\mathbf{W}_{i})$
5:end for
6:$\hat{y}$ = Fully Connected(Global Pooling($\mathbf{Y}_{k-1}$))
7:return $\hat{y}$
Lanczos Algorithm with Linear Constraints
For a given symmetric matrix $\mathbf{L}\in\mathbb{R}^{n\times n}$ and a randomly initialized vector $\bm{\nu}\in\mathbb{R}^{n}$, the $\kappa$-step Lanczos algorithm [27] computes an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n\times m}$ and a symmetric tridiagonal matrix $\mathbf{T}\in\mathbb{R}^{m\times m}$, such that $\mathbf{Q}^{\top}\mathbf{L}\mathbf{Q}=\mathbf{T}$. We represent $\mathbf{Q}_{N}=[\bm{q}_{1},\dots,\bm{q}_{N}]$ where the column vector $\bm{q}_{i}$ corresponds to the $i^{th}$ Lanczos vector. The matrices $\mathbf{B}\in\mathbb{R}^{m\times m}$ and $\mathbf{R}\in\mathbb{R}^{m\times m}$ represent the eigenvectors and eigenvalues of $\mathbf{T}$, respectively. By investigating the $j^{th}$ column of the system $\mathbf{L}\mathbf{Q}=\mathbf{Q}\mathbf{T}$ and rearranging terms, we obtain $\mathbf{L}\bm{q}_{j}=\beta_{j+1}\bm{q}_{j+1}+\beta_{j}\bm{q}_{j-1}+\alpha_{j}\bm{q}_{j}$.
Having the linear constraint changes the plain Lanczos algorithm by replacing $\bm{u}_{j}=\mathbf{L}\bm{q}_{j}-\beta_{j}\bm{q}_{j-1}$ with $\bm{u}_{j}=\bm{p}_{j}-\beta_{j}\bm{q}_{j-1}$ assuming the initial vector $\bm{\nu}$ is projected into the null space of the constraints [18]. Please note that the orthogonal projector $\mathbf{P}$ can be obtained through the QR decomposition of $\mathbf{C}$ when dealing with a dense constraint matrix $\mathbf{C}$. In situations where $\mathbf{C}$ is sparse and $\text{dim}(\mathcal{N}(\mathbf{C}^{\top}))\approx n$, the projector is given by $\mathbf{P}=\mathbf{I}-\mathbf{C}\mathbf{C}^{\dagger}$, with $\mathbf{C}^{\dagger}$ being the Moore-Penrose inverse of $\mathbf{C}$. Assuming $\mathbf{C}$ has full column rank, $\mathbf{C}^{\dagger}$ can be computed as $(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}$ [7]. If we project the initial vector $\bm{\nu}$ into null space of the constraint matrix $\bm{\nu}_{1}=\mathbf{P}\bm{\nu}\in\mathcal{N}(\mathbf{C}^{\top})$ and notice the mathematical equivalence between computing the smallest eigenvalue of the constraint $A_{p}=\mathbf{P}^{\top}\mathbf{L}\mathbf{P}$ and $\mathbf{L}$ then one step of the Lanczos algorithm with the linear constraints is $\beta_{j+1}\bm{q}_{j+1}=\mathbf{P}\mathbf{L}\mathbf{P}\bm{q}_{j}-\beta_{j}\mathbf{P}\bm{q}_{j-1}-\alpha_{j}\mathbf{P}\bm{q}_{j}=\mathbf{P}(\mathbf{L}\bm{q}_{j}-\beta_{j}\bm{q}_{j-1}-\alpha_{j}\bm{q}_{j})$. Algorithm 1 describes the steps of the Lanczos Algorithm with the Linear Constraints in detail.
This algorithm is structured into two main components: the ’outer loop’, which is a straightforward Lanczos algorithm iteration, and the ’inner loop’, which focuses on resolving the least squares problem expressed as $\mathbf{P}(\bm{b})=\underset{y\in\mathbb{R}^{l}}{\min}{\lVert\mathbf{C}\bm{y}-\bm{b}\rVert_{2}}$. Here, $\bm{y}$ is defined as $\mathbf{C}^{\dagger}\bm{b}$, and $\bm{b}$ is $\mathbf{L}\bm{q}_{j}$.
LLwLCNet
In this part, we detail our approach to computing the eigenvectors of the graph Laplacian matrix, ensuring they adhere to input graph constraints. We construct our eigenbasis by addressing a large, sparse, symmetric eigenvalue problem with homogeneous linear constraints. It requires minimizing
$\underset{\mathbf{C}{{}^{\top}}\bm{f}=0,\bm{f}\neq 0}{\min}\frac{\bm{f}{{}^{\top}}\mathbf{L}\bm{f}}{\bm{f}{{}^{\top}}\bm{f}},$ (1)
where $\mathbf{C}\in\mathbb{R}^{n\times l}$ with $n\gg l$ is the constraint matrix.
To address the problem outlined in Equation 1, we utilize the Lanczos algorithm with the linear constraints, detailed in Algorithm 1. However, differing from the iterative approach for the least square equation suggested in the Lanczos algorithm with the linear constraints to solve
$\mathbf{C}\bm{y}=\bm{b},$ (2)
we utilize the PyTorch framework [40] for our computations. This choice is due to our sparse and not overly large constraint matrix allowing for the direct QR factorization [2] within PyTorch, offering numerical stability and the capability for backpropagation. In the following, we describe constructing the constraint matrix $\mathbf{C}$, where we extract subgraphs first and derive the constraint matrix accordingly.
Constraint Matrix $\mathbf{C}$
The construction of the constraint matrix, denoted as $\mathbf{C}$, is based on subgraphs extracted from the input graph. We address the creation of $\mathbf{C}$ through two distinct approaches: induced subgraphs and link representation. For the induced subgraph case, we populate its corresponding column in $\mathbf{C}$ with the degrees of each node (and fill the remaining entries of the column with 00 for nodes not involved in the subgraph), ensuring that the derived eigenvectors fulfill the condition $\sum\bm{f}(x)d_{x}=0$. This condition ensures that the features learned using this eigenbasis reflect the imposed constraints. In the case of link representation, we design the constraint matrix to ensure that the aggregate of eigenvector differences across nodes equates to zero, as expressed by $\sum(\bm{f}(x)-\bm{f}(y))=0$. The column corresponding to the subgraph is constructed based on the equation. The degrees of the nodes involved in the constraint are entered into the column to satisfy the equation. Specifically, the degrees of nodes two-hop-away are negated, and the degrees of nodes one-hop-away are included only if they are connected to nodes two hops away. The remaining entries for nodes not involved are filled with 00 (see Figure 1 for an example). We build the other columns of the constraint matrix accordingly. A single matrix encompasses all constraints and each column representing a distinct constraint (${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{\sum}{}{}\bm{f}(x)d_{x}=0}$, $[{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}]$ ). Implementing these constraints and requiring the graph Laplacian matrix’s eigenvectors to comply with them allows for feature extraction aligned with the eigenbasis constraints. This enables nuanced differentiation between structures, such as $k$-regular graphs, as shown in Figure 1.
Addressing the eigenvalue problem with linear constraints yields a tridiagonal matrix, denoted as $\mathbf{T}$, and an orthogonal matrix, $\mathbf{Q}$. The decomposition of matrix $\mathbf{T}$ produces matrices $\mathbf{R}$ and $\mathbf{B}$. Here, $\mathbf{R}$ represents the Ritz eigenvalues, and $\mathbf{V}=\mathbf{Q}\mathbf{B}$ forms the eigenbasis that satisfies the constraints imposed by matrix $\mathbf{C}$. Forcing the eigenvectors to satisfy the constraints leads to having different eigenbasis for graphs where the MPNN returns the same features. This is exemplified in the case described in Figure 1, where two 2-regular graphs yield two different eigenbases.
Full Block
After determining the eigenbasis, we are ready to establish the full block of the Lanczos Layer with Linear Constraint (LLwLC). In this new eigenbasis, we develop spectral filters by applying a multilayer perceptron ($f$) to the eigenvalue matrix $\mathbf{R}$. With these learned filters, we reconstruct our basis and transform the graph signals $\mathbf{X}\in\mathbb{R}^{m\times n}$ into this basis to extract features that meet our specific constraints. $\hat{\mathbf{L}}$ is the graph Laplacian matrix computed from the low-rank approximation of the constrained eigenvalue problem. Each LLwLCNet block is
$\sigma(\mathbf{V}f(\mathbf{R})\mathbf{V}^{\top}\mathbf{X}\mathbf{W})=\sigma(\hat{\mathbf{L}}\mathbf{X}\mathbf{W}).$ (3)
Here, $\mathbf{W}\in\mathbb{R}^{n\times m}$ represents the learnable weight matrix, and $\sigma$ denotes the non-linearity applied in each block (ReLU in our experiments).
Full Architecture
As represented in Algorithm 2, we increase the number of blocks to deepen our architecture and capture more complex features. Each block reuses the initially computed eigenbasis and applies a multi-layer perceptron (MLP) to the eigenvalue matrix $\mathbf{R}$ to reconstruct its corresponding $\hat{\mathbf{L}}$. Our complete pipeline concludes by a global sort pooling [54] and a fully connected block in the last layer as depicted in Figure 1, used to predict link existence.
Lanczos Algorithm with Linear Constraint Convergence
In this section, we substantiate LLwLC eigenbasis’s convergence properties by conducting an error analysis on perturbations and referencing Greenbaum’s findings to demonstrate the existence of an exact Lanczos algorithm for any perturbed version. By establishing the upper bound for the Lanczos algorithm’s low-rank approximation, we affirm the convergence of our LLwLC eigenbasis.
Perturbation and Error Study
The accuracy of the linear least square problem using QR factorization depends on the precision of the QR factorization. As discussed by  Zhang, Baharlouei, and Wu [56], two types of accuracy errors are crucial in QR factorization when solving linear least square problems: The backward error for a matrix $\mathbf{Z}$ is defined as $\frac{\lVert\mathbf{Z}-\hat{\mathbf{Q}}\hat{\mathbf{R}}\rVert}{\lVert\mathbf{Z}\rVert}$ and the orthogonality error of $\hat{\mathbf{Q}}$ is measured by $\lVert\mathbf{I}-\hat{\mathbf{Q}}{{}^{\top}}\hat{\mathbf{Q}}\rVert$. Ideally, both numerical errors should be zero, but due to roundoff errors and the potential loss of orthogonality in the Gram-Schmidt QR process, the QR factorization might not be sufficiently accurate for solving the linear least square problem.
After examining the impact on accuracy, we analyze the theoretical gap between the exact Lanczos algorithm and its perturbed variant due to inexact QR factorization. The inexact QR factorization applied to solve 2 will impact the accuracy of both the Lanczos vectors and the tridiagonal matrices produced. Consequently, the computed tridiagonal matrix $\mathbf{T}_{j}$ is a perturbed version of the theoretical tridiagonal matrix, denoted as $\mathbf{T}^{\ast}_{j}$, that would be generated by an exact Lanczos iteration. This relationship can be expressed as $\mathbf{T}_{j}=\mathbf{T}^{\ast}_{j}+\mathbf{E}_{j}$, where $\mathbf{E}_{j}$ is the perturbation matrix after the $j^{th}$ step. The following theorem details the error bounds of the perturbed tridiagonal matrix in comparison to the theoretical exact solution of $\mathbf{T}$ after the $j^{th}$ step of the Lanczos algorithm.
Theorem 1.
Let $\mathcal{U}$ and $\tilde{\mathcal{U}}$ be the eigenspaces corresponding to the smallest eigenvalues $\lambda$ and $\tilde{\lambda}$ of the symmetric matrices $\mathbf{L}$ and $\tilde{\mathbf{L}}=\mathbf{L}+\mathbf{E}$, respectively. Then for any $\bm{u}\in\mathcal{U}$ and $\tilde{\bm{u}}\in\tilde{\mathcal{U}}$ with $\lVert\bm{u}\rVert_{2}=1$ and $\lVert\tilde{\bm{u}}\rVert_{2}=1$, we have $\tilde{\lambda}-\lambda\approx{\sum}_{i=1}^{j}\mathbf{E}_{j}(i,i)\bm{u}(i)^{2}+2{\sum}_{i=1}^{j-1}\mathbf{E}_{j}(i,i+1)\bm{u}(i)\bm{u}(i+1),$ where $\mathbf{E}_{j}(s,t)$ is the $(s,t)$ element of $\mathbf{E}_{j}$.
After exploring the theoretical gap between exact and perturbed Lanczos algorithms, we investigate Greenbaum’s result, which shows that each perturbed Lanczos corresponds to an exact version.
Greenbaum’s Results [19]
The tridiagonal matrix $\mathbf{T}_{j}$ generated at the end of the $j^{th}$ finite precision Lanczos process satisfying $\mathbf{L}\mathbf{Q}_{j}=\mathbf{Q}_{j}\mathbf{T}_{j}+\beta_{j+1}\bm{q}_{j+1}\bm{e}{{}^{\top}}_{j}+\mathbf{F}_{j},$ where $\bm{e}_{j}^{\top}$ is a vector with the $j^{th}$ component one and all the other components zero, $\mathbf{F}=(\bm{f}_{1},\dots,\bm{f}_{j})$ is the perturbation term with $\lVert\bm{f}_{j}\rVert_{2}\leq\epsilon\lVert\mathbf{L}\rVert_{2},\epsilon\ll 1$, is the same as that generated by an exact Lanczos process but with a different matrix $\tilde{\mathbf{L}}$. The matrices $\mathbf{L}$ and $\tilde{\mathbf{L}}$ are close in the sense that for any eigenvalue $\lambda(\tilde{\mathbf{L}})$ of $\tilde{\mathbf{L}}$, there is an eigenvalue $\lambda(\mathbf{L})$ of $\mathbf{L}$ such that $|\lambda(\tilde{\mathbf{L}})-\lambda({\mathbf{L}})|\leq\lVert\mathbf{F}_{j}\rVert_{2}.$ Therefore, in our case with the constant accuracy of the QR factorization, we can show $\mathbf{P}\mathbf{L}\mathbf{P}\tilde{\mathbf{Q}}_{j}=\tilde{\mathbf{Q}}_{j}\mathbf{T}_{j}+\beta_{j}\tilde{\bm{q}}_{j+1}\bm{e}_{j}^{\top}+\tilde{\mathbf{F}}_{j},$ where $\tilde{\mathbf{F}}_{j}=\mathbf{O}(\eta)$ with $\eta$ corresponds to the accuracy of the QR method.
Having established each perturbed Lanczos corresponds to an exact algorithm, we demonstrate the theorem below to bound the approximation error, as discussed in [29].
Theorem 2.
Let $\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$ be the eigendecomposition of an $n\times n$ symmetric matrix $\mathbf{L}$ with $\mathbf{\Lambda}_{i,i}=\lambda_{i},\lambda_{1}\geq\dots\geq\lambda_{n}$ and $\mathbf{U}=[\bm{u}_{1},\dots,\bm{u}_{n}]$. Let $\mathbf{\mathcal{U}}_{j}\equiv\text{span}\{\bm{u}_{1},\dots,\bm{u}_{j}\}$. Assume $\kappa$-step Lanczos algorithm starts with vector $\bm{\nu}$ and outputs the orthogonal $\mathbf{Q}\in\mathbb{R}^{n\times\kappa}$ and tridiagonal matrix $\mathbf{T}\in\mathbb{R}^{\kappa\times\kappa}$. For any $j$ with $1<j<n$ and $\kappa>j$, we have
$\lVert\mathbf{L}-{\mathbf{Q}}\mathbf{T}{\mathbf{Q}}^{\top}\rVert^{2}_{F}\\ \leq{\sum}_{i=1}^{j}\lambda^{2}_{i}\Biggl{(}\frac{\sin(\bm{\nu},\mathcal{U}_{i})\prod^{j-1}_{k=1}\frac{\lambda_{k}-\lambda_{N}}{\lambda_{k}-\lambda_{j}}}{\cos(\bm{\nu},\bm{u}_{i})T_{\kappa-i}(1+2\gamma_{i})}\Biggr{)}^{2}+{\sum}_{i=j+1}^{N}\lambda^{2}_{i},$
where $T_{\kappa-i}(x)$ is the Chebyshev Polynomial of degree $\kappa-i$ and $\gamma_{i}=(\lambda_{i}-\lambda_{i+1})/(\lambda_{i+1}-\lambda_{N})$.
Based on Greenbaum’s results [19], for our computed perturbed Lanczos algorithm, exists an exact Lanczos algorithm but for a different matrix. Based on 2, we also cognize the upper bound of the low-rank approximator of the Lanczos algorithm. Thus, the perturbed Lanczos algorithm, caused by the inaccuracy of the QR method for solving the least square equation, converges to the upper bound of the low-rank approximation of the matrix of the exact Lanczos algorithm.
Subgraph Extraction Policy
A subgraph selection policy is a function $\pi\colon\mathcal{G}\rightarrow\mathbb{P}(\mathcal{G})$ assigning to a graph a subset of its subgraphs [6]. Here, $\mathcal{G}$ is the set of all graphs with $n$ nodes or less and $\mathbb{P}(\mathcal{G})$ its power set. Although any linear constraint in the input graph satisfying full rank assumption can be encoded in $\mathbf{C}$, we propose the following subgraph extraction policies.
Neumann Eigenvalue
The Neumann eigenvalue [15] is
$\lambda_{S}=\inf_{\bm{f}\neq 0}\frac{{\sum}_{x\in S}\bm{f}(x)\mathbf{L}\bm{f}(x)}{{\sum}_{x\in S}\bm{f}^{2}(x)d_{x}},\text{ subject to}$
$\sum_{y\in S,x\in\delta S,y\sim x}(\bm{f}(x)-\bm{f}(y))=0\quad\text{and}\quad\sum_{x\in S}\bm{f}(x)d_{x}=0.$
The function $\bm{f}\colon S\cup\delta S\rightarrow\mathbb{R}$ represents the Neumann eigenvector satisfying the Neumann conditions. The vertex boundary, $\delta S$, of an induced subgraph consists of all vertices not in $S$ but adjacent to at least one vertex in $S$. Specifically, the first constraint encodes the link representation ($\sum_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y}\in S,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y}\sim{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}x}}(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{f}(x)})-({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\bm{f}(y)})=0$, $[{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}0},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1},{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}-1},{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}]$ ; building on previous link prediction research, we consider nodes that are two hops away from the query nodes, where $S$ represents the one-hop-away nodes, and $\delta S$ denotes the boundary nodes between one-hop and two-hop-away nodes). This can be reformulated as:
$\min_{\bm{f}\in\mathbb{R}^{n}}\bm{f}^{\top}\mathbf{L}\bm{f},\text{ subject to }\lVert\bm{f}\rVert=1\text{ and }\mathbf{C}^{\top}\bm{f}=0,$ (4)
where $\mathbf{L}\in\mathbb{R}^{n\times n}$ is a symmetric and large sparse matrix, and $\mathbf{C}\in\mathbb{R}^{n\times l}$ (with $n\gg l$) is also large, sparse, and of full column rank. The time complexity involved in extracting subgraphs depends on the product of the maximum degree of nodes and the count of nodes within nodes in the boundary. When we enforce that the eigenvectors satisfy the constraints related to induced subgraphs and link representation, we ensure that the corresponding features adhere to these constraints.
Proposition 1.
Applying Neumann eigenvalue constraints to the eigenbasis results in features that exhibit greater expressivity than MPNNs. Besides addressing the node automorphism problem [44], these enhanced features enable the distinction of specific $k$-regular graphs from each other, thereby significantly enhancing expressivity in GNNs (Proof in Appendix).
Constraints $\mathbf{C}$
The second subgraph extraction policy we propose is based on vertex-deleted subgraphs. The sufficient conditions under which LLwLC can solve graph isomorphism entails that LLwLC is a universal approximator of functions defined on graphs [14]. Given that we can encode any subgraph into our eigenbasis, we can examine whether a specific substructure collection can completely characterize each graph. By the reconstruction conjecture [45], we assume we can reconstruct the graph if we have all the $n-1$ vertex-deleted subgraphs. According to [26], exponentially many graphs are determined by their spectrum. Thus, it is believable that the same conjecture holds for the eigenbasis of graph Laplacian matrices of exponentially many graphs.
In extracting vertex-deleted subgraphs from a dense graph, the worst-case time complexity is $\mathcal{O}(n^{3})$. This complexity is due to the high number of edges in dense graphs, leading to increased computational load. For sparse graphs with fewer edges, the complexity is reduced by a factor of $\mathcal{O}(n)$. Our empirical results show that applying a few constraints significantly improves performance, making the practical overhead align linearly with the number of nodes. This approach is similar to the marker-based method [38], as both consider slightly perturbed input graphs.
Utilizing vertex-deleted subgraphs allows us to distinguish between graphs that are not distinguishable with the 2-WL method. For instance, although the $4\times 4$ rook and Shirkhande graphs are indistinguishable with 2-WL, their corresponding constraint matrices—constructed from the vertex-deleted subgraphs—have different spectra and thus different constraint matrices. These differences in spectra lead to distinct eigenbases, enabling us to differentiate between these graphs with our novel LLwLC.
Furthermore, we conjecture that if the reconstruction conjecture holds [45] and the assumption that almost all graphs are reconstructable by their spectrum [26] is satisfied and the substructure collection contains all graphs of size $k=n-1$ with the form of ${\sum}_{x\in S}{}\bm{f}(x)d_{x}=0$ (for every vertex-deleted subgraph $S$), then it is believable and we conjecture that LLwLCNet can distinguish almost all non-isomorphic graphs of size $n$ and is therefore universal.
Stochastic Constraints $\mathbf{C}$
In our study, we empirically discovered that effective outcomes can be attained with a limited number of constraints, such as Neumann constraints, or in another ablation study involving only ten vertex-deleted subgraphs. This observation aligns with the findings of Bollobás [10], who demonstrated that almost all graphs can be reconstructed using only three vertex-deleted subgraphs. Given the prohibitive cost of considering all $n$ vertex-deleted subgraphs, our approach, inspired by  [6], involves selecting a subset of constraints for consideration. During each epoch, we stochastically choose $k$ vertex-deleted subgraphs for analysis.
Experiments
To demonstrate the effectiveness of our method in addressing node automorphism and leveraging subgraphs, we conducted experiments focused on the link prediction task. We compared the performance of LLwLCNet against traditional heuristics (CN [5], RA [57], AA [1]), vanilla GNNs (GCN [25], SAGE [21]), GNNs modifying the input graph of MPNNs (SEAL [53], NBFNet [58]), and GNNs with manual features as pairwise representations (Neo-GNN [51], BUDDY [13]). Baseline results are from Chamberlain et al. [13]. Our evaluation includes five link prediction benchmarks: Cora, Citeseer, Pubmed [49] (Planetoid datasets), OGBL-Collab, and OGBL-Vessel [24]. Dataset statistics are shown in Appendix. Baseline results for OGBL-Collab and OGBL-Vessel are from the OGB leaderboard.
Setup
In our experiments, we used a learning rate of 0.001 for 20 training epochs. The model has two 32-channel MLP layers, each with ReLU non-linearities and dropout. We cap the eigenpairs at 10, padding missing pairs with zeros. Both PyTorch [40] and PyTorch Geometric [17] were used for implementation. The binary cross entropy loss measured training loss. Following SEAL, we altered 10% of links for test data and used the remaining 90% for training.
Link Prediction Results
While the LLwLC framework theoretically applies to various problem domains, our experiments focus on the LP tasks to demonstrate its effectiveness in addressing node automorphism, considering substructures, and thus improving expressiveness. Consistent with prior studies [13], considering two-hop nodes is sufficient for effective LP task.
As depicted in Table 1, LLwLC stands out as a robust framework for link prediction, consistently delivering impressive performance on link prediction benchmarks with metrics given in the first row. With only 0.02M, LLwLC outperforms previous models on the Planetoid dataset, highlighting the significance of encoding subgraph structures and selective subsets of node relations for superior link prediction results. Our model also achieves state-of-the-art performance on the OGBL-Collab dataset with just 0.02M, compared to BUDDY (1.10M) and SEAL (0.50M). Increasing blocks to 0.03M yields a 67.50% HR@50 score. On OGBL-Vessel, we achieve competitive results with just 0.019M and 10% of the training data, demonstrating LLwLC’s effectiveness in a lightweight architecture.
Table 1: Results on LP benchmarks; LLwLCNet w. Neumann Constraints. The colors denote the best and second-best models. LLwLCNet trained with $10\%$ of the VESSEL dataset.
Cora Citeseer Pubmed Collab Vessel
Method HR@100 HR@100 HR@100 HR@50 roc-auc
CN 33.92 29.79 23.13 56.44 48.49
AA 39.85 35.19 27.38 64.35 48.49
RA 41.07 33.56 27.03 64.00 n.a.
GCN [25] 66.79 67.08 53.02 44.75 43.53
SAGE [22] 55.02 57.01 39.66 48.10 49.89
Neo-GNN [51] 80.42 84.67 73.93 57.52 n.a.
SEAL [53] 81.71 83.89 75.54 64.74 80.50
NBFnet [58] 71.65 74.07 58.73 OOM n.a.
Surel+ [50] N.A. N.A. N.A. 64.10 85.73
BUDDY [13] 88.00 92.93 74.10 65.94 55.14
LLwLCNet 91.44 93.40 83.10 66.86 81.60
# Params. $0.019$M $0.018$M $0.024$M $0.026$M $0.036$M
Table 2: Time Complexity Comparison
(a) LP Tasks’ Time Complexity: Number of Nodes (n), Edges (E), Dimensionality of Node Features (d), Propagation Hops (l), Sketch Size in BUDDY (h) [13], Eigenvectors ($\kappa$), and Constraints ($k$).
Time Complexity SEAL BUDDY NBFNet LLwLC
Preprocessing $\mathcal{O}(1)$ $\mathcal{O}(lE(d+h))$ $\mathcal{O}(1)$ $\mathcal{O}(1)$
Training (1 link) $\mathcal{O}(Ed^{2})$ $\mathcal{O}(l^{2}h+ld^{2})$ $\mathcal{O}(Ed+nd^{2})$ $\mathcal{O}(\kappa E+k^{2}n)$
Inference $\mathcal{O}(Ed^{2})$ $\mathcal{O}(l^{2}h+ld^{2})$ $\mathcal{O}(Ed+nd^{2})$ $\mathcal{O}(\kappa E+k^{2}n)$
(b) Expressivity Models’ Time Complexity.
Time Complexity $k$-IGN $k$-GNN LLwLC GSN (3-WL)
Preprocessing - - $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{k})$
Dense $\mathcal{O}(n^{k})$ $\mathcal{O}(n^{k})$ $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$
Sparse $\mathcal{O}(n^{k})$ $\mathcal{O}(n^{k})$ $\mathcal{O}(n)$ $\mathcal{O}(n)$
Time Complexity
The time complexity of our method is $\mathcal{O}(\kappa E)+\mathcal{O}(k^{2}n)+\mathcal{O}(k^{3})$ for the outer loop (Lanczos algorithm), the QR factorization, and the pseudo-inverse of $k\times k$ matrix, respectively, where $\kappa\ll n$ is the number of computed eigenvectors and $k\ll n$ is the number of linear constraints. Table 2(a) compares the time complexities of LLwLCNet and other link prediction methods.
Comparing the LLwLCNet model’s time complexity with $k$-WL expressive models provides valuable insights. It is conjectured that LLwLCNet is a universal approximator with a worst-case time complexity of $\mathcal{O}(n^{3})$ (where $k=n-1$). Aligned with [5] research and supported by our empirical experiments, it has been established that graph reconstruction can be effectively achieved with just a few constraints. This leads to time complexity linear to the number of nodes. Contrastingly, established $k$-WL expressive models, such as $k$-IGNs [30] and $k$-GNNs [34], are known to have a higher time complexity of $\mathcal{O}(n^{k})$.
Furthermore, we compare LLwLCNet’s time complexity with GSN [11]. It explores substructures by counting specific substructures. However, it relies on prior knowledge and inductive bias for task-specific substructure selection and deals with the subgraph isomorphism. In GSN, the preprocessing step, in general, is $\mathcal{O}(n^{k})$ for a generic substructure of size $k$. In contrast, our experimental results demonstrate that LLwLCNet can capture graph properties without requiring inductive bias. Comparisons are in Table 2(b).
Ablation Studies
To demonstrate the impact of the applied constraints, we compare the eigenbasis estimation using the ground truth Laplacian matrix $\mathbf{L}$ (with edges provided) against the approximation of the eigenbasis using the ground truth Laplacian matrix $\mathbf{L}$ alongside Neumann constraints $\mathbf{C}$. This comparison is conducted on three benchmark datasets, as detailed in Table 3(a). We notice that including Neumann constraints, $\mathbf{C}$, markedly enhances the results.
In our extended research, we explore the impact of adding additional constraints to our framework. As detailed in Table 3(b), we explore the influence of integrating an increased number of constraints using vertex-deleted subgraphs. Our observations reveal that implementing merely ten constraints from vertex-deleted subgraphs yields state-of-the-art enhancements in benchmark datasets. Notably, within the OGBL-Collab dataset, the constraints significantly boost performance metrics (e.g., HR@50 increases from $42.83$ in LanczosNet, which does not apply linear constraints to the input graph, to $69.40$ with the incorporation of ten vertex-deleted subgraph constraints). Similar improvements are evident in the Cora dataset, with HR@100 rising from $90.80$ to $93.10$, and PubMed shows similar improved results over the baseline.Our results are consistent with the theoretical framework. It demonstrates that incorporating $n-1$ vertex-deleted subgraphs as linear constraints leads to a universal approximator function. Significantly, it aligns with [10], suggesting nearly all graphs can be reconstructed using three vertex-deleted subgraphs.
Table 3: Subgraph Constraint Impact. Best in red.
(a) Results with/without Neumann constraints (with $\mathbf{L}$).
Cora Citeseer PubMed
Method AUC AUC AUC
LanczosNet 94.5% 96.5% 97.2%
LLwLC (w. $\mathbf{L}$ $\mathbf{C}$) 97.0% 98.1% 98.3%
(b) Results with different subgraph constraints.
Collab Cora PubMed
Method HR@50 HR@100 HR@100
LanczosNet [29] 42.58 90.80 77.18
LLwLCNet w. Neumann Constraints 66.86 91.44 83.10
LLwLCNet w. 10 Constraints 69.40 93.10 82.28
# Params. 0.026M 0.019M 0.021M
Wall-time and Reduction Dataset
Our study has shown an enhancement in model expressivity, leading to a decrease in the need for parameters. This results in less training data being required. As shown in Table 4, on the PubMed dataset, our model surpasses BUDDY [13] and SEAL [53] using merel $10\%$ of the training data, which results in a training speed that is 20 times faster than SEAL. Furthermore, it delivers performance on par with BUDDY while using just $1\%$ of the training data within the same training duration and without requiring any preprocessing steps. This enhanced efficiency reflects a more efficient learning mechanism and significantly reduces training duration. Regarding the OGBL-Vessel dataset, our approach outperforms SEAL [53], demonstrating a 10-fold increase in training speed. It achieves similar outcomes at a speed 90 times faster using only 1% of the training data. These experiments were all carried out on a GeForce GT1030 GPU (CUDA 11.6, PyTorch 1.13).
Table 4: Wall-time comparisons with LP models. Training time is one epoch.
Dataset SEAL BUDDY GCN LLwLC LLwLC LLwLC
Pre-train(s) 0 5 0 0 0 0
Train (s) 81 1 66 7 4 1
PubMed Reduction 100 100 100 10 5 1
Accuracy 75.54 74.10 53.02 80.15 78.20 70.84
# Params. 0.486M 1.565M 0.052M 0.024M 0.024M 0.024M
Train (s) 12600 N.A. 12600 1200 700 140
OGBL-Vessel Reduction 100 100 100 10 5 1
Accuracy 80.50 55.14 43.53 81.60 79.24 78.72
# Params. 0.042M N.A. 0.035M 0.036M 0.036M 0.036M
Related Work
MPNN Expressivity
The expressivity of GNNs is typically expressed in terms of their ability to distinguish non-isomorphic graphs. As no polynomial-time algorithm for solving the graph isomorphism problem is known, developing GNNs that are both expressive and efficient poses a major challenge. Xu et al. [48] found that the expressivity of MPNNs is limited to that of the 1-WL test. This limitation is crucial in real-world applications, as the 1-WL test cannot distinguish certain graph structures, such as isomorphism in attributed regular graphs, measuring node distances, and counting cycles [28]. Recent studies have addressed these limitations with four approaches: adding random attributes to nodes [41], using deterministic positional features [53], developing higher-order GNNs to surpass the 1-WL test’s expressivity limits [31], and subgraph GNNs applying markings, such as the node-deletion approach of ESAN [6]. Please refer to the recent survey [33] for a comprehensive overview of these techniques. Our approach aligns with the direction of subgraph GNNs.
Subgraph GNNs for Link Prediction
SEAL enhances WLNM [52], the first subgraph-based LP, by using graph convolutional layers and encoding positional features. SEAL demonstrates that information within two-hop subgraphs is sufficient, aligning with classical methods. Neo-GNN [51] and BUDDY [13] decouple pairwise representation from node representation learning to reduce computational overhead but may oversimplify pairwise representations.
Conclusion
In this work, we introduced the Learnable Lanczos algorithm with Linear Constraints (LLwLC), a novel method designed to enhance the expressivity of Graph Neural Networks (GNNs). Through the incorporation of two novel subgraph extraction strategies, we managed to construct a lightweight architecture that minimizes reliance on extensive training datasets. Empirical results show that our method significantly improves performance in link prediction tasks across various benchmark datasets. Notably, the LLwLC achieved 20$\times$ and 10$\times$ speedup, requiring only 5% and 10% data from the PubMed and OGBL-Vessel datasets respectively, compared to the state-of-the-art methods. These findings underscore the practical utility and theoretical advancement of our method, illustrating the LLwLC’s potential as a universal approximator and its ability to differentiate between $k$-regular graphs. The advancements made with the LLwLC not only represent a significant contribution to the field but also set a promising direction for future exploration and development in the realm of GNNs. As a further future work, we believe investigating the impact of learning linear constraints between nodes and edges within the input graph and encoding them to the eigenbasis of the graph Laplacian matrix, can promise further advancements in this field.
Acknowledgment This research was funded by the Austrian Research Promotion Agency (FFG) under project no. 874065. The author sincerely appreciates the valuable feedback from Mohsen Fayyaz, Joshua Erde, and András Papp.